英語 (English)

主題別英語 (Thematic English) (工 ((建 A)1 年))

バロッグスザンネ · 非常勤講師 / 全学共通教育センター 1 単位後期水 5 ・ 6

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 English has become the lingua franca in today’s global society. As globalization brings the world closer together, the need for global citizens to be able to communicate clearly in English is becoming more urgent. The aim of this course is to give students the opportunity to practice speaking English with an emphasis on correct pronunciation, rhythm and intonation. Listening to authentic native speaker pronunciation (through DVD and other media) will be an important part of the course. Practice in the four skills reading, writing, listening and speaking will be provided during class and through homework.

【授業の概要】Students will be expected to take part in a number of role-plays and memorize short skits based on movie scenes. Emphasis will be placed on producing good English pronunciation with natural rhythm and intonation. The following schedule is tentative and shows only general topics for each class. It may be modified depending on students’ level and interest.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】

1. Students will gain confidence in their pronunciation and speaking skills.

2. Students will have improved listening comprehension and a more thorough understanding of the English phonetic system and other prosodic features of English.

【授業の計画】 1. Introductions

2. Background to The Sound of Music

3. Listening 1: Maria meets the children

4. Listening 2: In Maria’s room

5. Listening 3: Song: Do Re Mi

6. Listening 4: The Captain returns

7. Listening 5: Maria returns to the abbey

8. Listening 6: Climb every mountain 9. Listening Test

10. Listening 7: Maria and the Captain

11. Listening 8: Maria and Liesl

12. Listening 9: Escape from Austria

13. The ‘real’ Maria and her life

14. Listening Test 2

15. Meeting the children 40 years on

16. Consolidation

【教科書】 Handouts will be prepared by the instructor and no textbook is required. 【参考書等】 [参考資料] 【成績評価の方法】 Students will be assessed on class participation, listening tests and written assignments. Tentative coursework evaluation: Attendance: 30%, Listening Tests: 40%, Written Assignments: 30%

【再試験の有無】 Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. 【受講者へのメッセージ】 [注意]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221175 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ バロッグ .

English

Thematic English (工 ((建 A)1 年))

Susanne Balogh · Part-time Lecturer / Center for General Education 1 unit 後期水 5 ・ 6

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ English has become the lingua franca in today’s global society. As globalization brings the world closer together, the need for global citizens to be able to communicate clearly in English is becoming more urgent. The aim of this course is to give students the opportunity to practice speaking English with an emphasis on correct pronunciation, rhythm and intonation. Listening to authentic native speaker pronunciation (through DVD and other media) will be an important part of the course. Practice in the four skills reading, writing, listening and speaking will be provided during class and through homework.

Outline⟩ Students will be expected to take part in a number of role-plays and memorize short skits based on movie scenes. Emphasis will be placed on producing good English pronunciation with natural rhythm and intonation. The following schedule is tentative and shows only general topics for each class. It may be modified depending on students’ level and interest.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩

1. Students will gain confidence in their pronunciation and speaking skills.

2. Students will have improved listening comprehension and a more thorough understanding of the English phonetic system and other prosodic features of English.

Schedule⟩

1. Introductions

2. Background to The Sound of Music

3. Listening 1: Maria meets the children

4. Listening 2: In Maria’s room

5. Listening 3: Song: Do Re Mi

6. Listening 4: The Captain returns

7. Listening 5: Maria returns to the abbey

8. Listening 6: Climb every mountain 9. Listening Test

10. Listening 7: Maria and the Captain

11. Listening 8: Maria and Liesl

12. Listening 9: Escape from Austria

13. The ‘real’ Maria and her life

14. Listening Test 2

15. Meeting the children 40 years on

16. Consolidation

Textbook⟩ Handouts will be prepared by the instructor and no textbook is required. Reference⟩ [参考資料] Evaluation Criteria⟩ Students will be assessed on class participation, listening tests and written assignments. Tentative coursework evaluation: Attendance: 30%, Listening Tests: 40%, Written Assignments: 30%

Re-evaluation⟩ Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. Message⟩ [注意]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221175 Contact (Office-Hour, Room, E-mail)⟩

⇒ Balogh .